# A DNN Biophysics Model with Topological and Electrostatic Features

### v2

@@ Abstract
In this project, we provide a deep-learning neural network (DNN) based biophysics model to predict protein properties. The model uses multi-scale and uniform topological and electrostatic features generated using protein structural information and force field. The topological features are generated using the element specified persistent homology (ESPH) on a selection of heavy atoms or carbon atoms which contribute significantly in the protein biophysical properties. The electrostatic features are fast computed using a Cartesian treecode, modulated with treecode parameters for accuracy-cost tradeoff. These features are uniform in number for proteins with various sizes thus the broadly available protein structure database can be used to train the network. These features are also multi-scale thus the resolution and computational cost can be balanced by the users at the consideration of available computing powers. The machine learning simulations on two data sets with 4000+ and 17000+ protein structures showed the efficiency and fidelity of these features in representing the protein structure and force field for the predication of their biophysical properties such as electrostatic Coulombic energies and solvation energies. The feature generation algorithms have the potential as general tools in assisting machine learning based biophysical properties and function prediction for the broad biomolecules whose structures are obtained from both theoretical computing and experiments.
keywords:
Electrostatics; Poisson-Boltzmann; treecode; Deep Neural Network; Persistent Homology
@@ 1 Introduction
One of the overarching themes of biology is that structure determines function, which becomes more significant when we explore the biological world at the molecular level. With the advancement and availability of structure determination techniques such as X-ray crystallography [1, 2], NMR spectroscopy [3], cryo-electron microscopy [4], etc., there are rapidly growing 3D structures of protein, nucleic acids, and complex assemblies, etc. on publicly accessible repository such as Protein Data Bank (PDB) [5], GenBank [6], etc. These structure determination techniques have strengths and limitations that make them suitable for different types research studies. For examples, X-ray crystallography, due to the requirement of crystallization, is best for static high-resolution structures. NMR is valuable for studying dynamics in solution, particularly for the studies of smaller proteins and complexes. Cryo-electron microscopy which becomes more and more popular recently, excels in visualizing large, complex, and flexible biomolecular systems. The combination and integration of these methods due to their complementary natures help to reveal critical biomolecular functions embedded in the structures.
@@
In addition to experiments, computing approaches are also available for the prediction of protein structures and properties. The main stream of Machine Learning based protein prediction focuses on using amino acid sequence to predict protein structure or structure-related properties, resulting in the trending protein language models (PLM) [7, 8], which is a transformer-based model inspired by Natural Language Processing (NLP) with wide range of applications such as protein pKa prediction [9], protein structure prediction [10], protein design [11], protein evolution [12], etc.
Recently, the appearance of AlphaFold developed by Google DeepMind revolutionarily reshaped the structures determination techniques. Based on known structures, AlphaFold learned to predict protein’s structures from sequence with very high accuracy as shown by their victory by a large margin in Critical Assessment of Techniques for Protein Structure Prediction (CASP 14-16), AlphaFold Protein Structures Database now has over 200 million protein structure predictions [13].
@@
Both PLM and AlphaFold take protein sequences as inputs to predict structures or embeddings which keep important features, without touching the prediction of protein functions from protein structures. The machine learning methods potentially can revolutionize the traditional structure-based molecular modeling, which uses brownian dynamics, molecular dynamics, Monte Carlo simulations, etc. to predict protein properties and functions from protein structures. However, the development of ML methods in protein property predictions is hindered by the difficulty to represent protein structures and force fields using reliable features. To this end, the structure data, which is largely variant from protein to protein, needs to be converted to uniform size of features in representing structures to be loaded into the machine learning models such as regression models, random forest, deep neural network etc. Successful cases are the use of graph convolutional network (GCN) for predicting protein functions by features extracted from a protein language model and protein structure [14], protein-DNA binding specificity prediction using a wide range of features from geometry [15, 16], protein-ligand binding affinity prediction using topological features, [17, 18].
@@
In this paper, we provide a novel way to generate electrostatic features. Combining electrostatic features with the topological features generated using a modification of the persistent homology approach used in [19], we represent the protein structures and associated force field in a uniform and multiscale fashion. With these features, we train a deep-learning neural network (DNN) to predict protein Coulombic energy and electrostatic solvation energy efficiently and accurately. Here the Solvation energies as labels are generated by solving the Poisson-Boltzmann (PB) model using a second-order Matched Interface and Boundary (MIB) method based PB solver [20]. There are recent research using Physics Informed Neural Network (PINN) to solve the Poisson-Boltzmann model [21, 22] while the present work is rather different as a Data-driven method with uniform and multi-scale features. In addition to these particular application in predicting Coulombic energy and electrostatic solvation energy, the feature generation algorithm we have developed potentially has wider application for other machine learning models as far as the protein structures and force field are required to be represented uniformly for proteins with various sizes.
@@
This manuscript is constructed in the following way. After this section of introduction is the section of models and algorithms, followed by the section of results of simulation. The manuscript is ended with a conclusion section.
2 Models and Algorithms
@@
We introduce the models and algorithms involved in this work. For protein’s structure-property relation we focus on the atomic description and the electrostatic interactions. The protein properties involved in this paper are Coulombic energy $E_{\text{coul}}$ and electrostatic solvation energy $E_{\text{solv}}$. For a protein, the Coulombic energy $E_{\text{coul}}$ is the energy it takes to displace atomic charges of the protein from infinity to their present positions (i.e. atomic centers from the PDB file) while the electrostatic solvation energy $E_{\text{solv}}$ is the energy it takes to solvate the protein into the solvent (e.g. water). Formal quantitative definition of these two protein properties will be given with mathematical models. For the protein-solvent interaction, we use the implicit solvent model in which water is represented as a continuum for the reduced computational cost. In a comparison perspective, we introduce the two most popular implicit solvation models, namely the Poisson-Boltzmann model and the Generalized Born’s model, which are used to generate the labels or features. Following that, we introduce the persistent homology and how we use it to produce the topological features. Finally we introduce how to use the Cartesian treecode, whose modification evolves into the algorithm to generate the electrostatic features.
@@ 2.1 The Poisson-Boltzmann model
Figure 1 depicts the popular implicit solvent models. In Fig.  1(a) a protein is represented by a collection of $N_{c}$ spherical atoms with centered partial charges. The molecular surface $\Gamma$ (also known as the solvent excluded surface) is defined by the trace of a water molecule represented by a red sphere rolling on contacting with the protein atoms. The Poisson-Boltzmann model is shown in Fig. 1(b), where the molecular surface $\Gamma$ divides the entire computational domain $\Omega$ into the protein domain $\Omega_{1}$ with dielectric constant $\varepsilon_{1}$ and atomic charges $q_{k}$ located at ${\bf x}_{k},k=1\,{:}\,N_{c}$, and the solvent domain $\Omega_{2}$ with dielectric constant $\varepsilon_{2}$ and dissolved salt ions  [23, 24]. Assuming a Boltzmann distribution for the ion concentration, and considering the case of two ion species with equal and opposite charges (e.g. Na+, Cl-), in the limit of weak electrostatic potential one obtains the linearized Poisson-Boltzmann (PB) model [25],
$\varepsilon_{1}\nabla^{2}\phi_{1}(\textbf{x})$ $=-\sum_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k}),\quad{\bf x}\in\Omega_{1},$ (1a)
$\varepsilon_{2}\nabla^{2}\phi_{2}(\textbf{x})$ $=\kappa^{2}\phi_{2}(\textbf{x}),\quad{\bf x}\in\Omega_{2},$ (1b)
@@
Figure 1: Implicit solvent models. (a) The Molecular Surface $\Gamma$: the trace of the solvent probe (shown in red) when it is rolled on contacting the spherical atoms of the protein (shown in green with a centered charge); (b) The Poisson–Boltzmann Model: two domains $\Omega_{1}$ (dielectric constant $\epsilon_{1}$ with partial charges as weighted summation of delta functions) and $\Omega_{2}$ (dielectric constant $\epsilon_{2}$ with mobile ions modeled by Boltzmann distribution) are separated by the molecular surface $\Gamma$; (3) The Generalized Born Model: the protein is represented as a collection of $N_{c}$ spherical atoms centered at ${\bf x}_{k}$ with charge $q_{k}$ and Born Radius $R_{k}$ (only the $k$th atom is shown).
The PB model governs the electrostatic potential $\phi$ in the entire space. Theoretically after $\phi$ is obtained, its gradient will produce electrostatic field while its integral will generate potential energy. However, there are many challenging issues on properly obtaining the field and energy (e.g. definition of field on molecular surface $\Gamma$ [26]). Our attention for this project is on the energy side as described below. The electrostatic free energy is given as
@@
where $\rho({\bf x})=\sum\limits_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k})$ is the charge density as a sum of partial charges weighted delta function and the $E_{\text{solv}}=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}\phi_{\text{reac}}({\bf x}_{k})$ term is the electrostatic solvation energy. The $\phi_{\text{reac}}({\bf x}_{k})$ is the reaction potential at ${\bf x}_{k}$ as the remaining component when Coulomb potential $\phi_{\text{coul}}({\bf x}_{k})=\sum\limits_{j=1,j\neq k}^{N_{c}}\frac{q_{j}}{\epsilon_{1}|{\bf x}_{k}-{\bf x}_{j}|}$ is taken away from the total electrostatic potential $\phi({\bf x}_{k})$.
Solving the PB model numerically by grid-based methods is challenging because (1) the protein is represented by singular point charges, (2) the molecular surface is geometrically complex, (3) the dielectric constant is discontinuous across the surface, (4) the domain is unbounded. To overcome these numerical difficulties, several finite difference interface methods have been developed. By assuming that the interface is aligned with a mesh line, a jump condition capture scheme has been developed in [27]. Based on a Cartesian grid, the immersed interface method (IIM) [28] has been applied to solve the PB equation [29, 30], in which the jump conditions can be rigorously enforced based on Taylor expansions. For the purpose of dealing with arbitrarily shaped dielectric interfaces based on a simple Cartesian grid, a matched interface and boundary (MIB) PB solver [31, 32, 33, 20, 34] has been developed through rigorous treatments of geometrical and charge singularities. Alternatively, boundary element methods (BEM) for the PB model were developed later [35, 36, 37, 38, 39, 40, 41, 42] with several inherent advantages, (1) only the molecular surface is discretized rather than the entire solute/solvent volume, (2) the atomic charges are treated analytically, (3) the interface conditions are accurately enforced, (4) the far-field boundary condition is imposed analytically, (5) the matrix-vector product in iterative methods can be accelerated by fast algorithms [43, 44, 45, 46]. In this paper, we computed the solvation energies using the MIBPB solver [20] since accuracy is of critical importance for the generation of labels ($E_{\text{solv}}$) that will be used to train the neural network.
@@ 2.2 Generalized Born’s Model
If ion effects are not considered, the solvation energy and the reaction potential can also be efficiently computed by using the generalized Born’s (GB) model, an approximation to the Poisson’s model. The GB model has significantly reduced computational cost as opposed to the PB or Poisson models, whose solutions require solving a 3D partial differential equation. Although we did not use the GB model to produce any numerical results for this project, we describe the GB model for its relationship with the PB model and its potential usage in electrostatic interactions. In fact, the GB model can calculate the electrostatic solvation energy which can be used either as the label or as a feature [47, 48] in our machine learning framework. It can also produce the reaction potentials at the atomic center, which can be used to generate the second level of electrostatic features as shown in Fig. 3, which will be explained within that content.
To derive the GB model, we first adopt some classic results for electrostatics from Jackson’s textbook [49] (sec. 1.11). Using the analogue from discrete point charge distribution for continuous charge density distribution, the electrostatic free energy of a charged system with density distribution $\rho({\bf r})$ takes the form
$E_{\text{free}}=\frac{1}{2}\int\rho({\bf r})\phi({\bf r})d{\bf r}=\frac{1}{8\pi}\int{\bf E}({\bf r})\cdot{\bf D}({\bf r})d{\bf r}$
by using the Poisson equation $\nabla^{2}\phi=\rho({\bf r})/\epsilon$, the electric field ${\bf E}=-\nabla\phi$, and electric displacement ${\bf D}=\epsilon{\bf E}$.
Now consider assembling a charge $q_{i}$ to the center of a sphere of radius $a_{i}$ and the center is located at the origin. The sphere separates the domain with $\epsilon_{1}$ for $r<a_{i}$ and $\epsilon_{2}$ for $r>a_{i}$. The assembly takes the energy
$E_{\text{free,}i}=\frac{1}{8\pi}\int\frac{{\bf D}\cdot{\bf D}}{\epsilon}d{\bf r}\approx\frac{1}{8\pi}\int_{r<a_{i}}\frac{q_{i}^{2}}{r^{4}\epsilon_{1}}d{\bf r}+\frac{1}{8\pi}\int_{r>a_{i}}\frac{q_{i}^{2}}{r^{4}\epsilon_{2}}d{\bf r}$ (3)
where the Coulomb field approximation ${\bf D}=\displaystyle\frac{q_{i}{\bf r}}{r^{3}}$ is used.
From a slightly different perspective as in Eq. (2), the electrostatic solvation energy is the energy difference when $\epsilon$ for $r>a_{i}$ changes from $\epsilon_{1}$ (unsolvated) to $\epsilon_{2}$ (solvated) given as
$E_{\text{solv},i}=\frac{1}{8\pi}(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\int_{r>a_{i}}\frac{q_{i}^{2}}{r^{4}}d{\bf r}=(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\frac{q_{i}^{2}}{2a_{i}}.$ (4)
This result is consistent with the Poisson-Boltzmann model of a solvated spherical cavity with centered charges as summarized in [50], which is a special case from Kirkwood’s derivation of a series of spherical harmonics for a spherical cavity containing arbitrary multiple charges [51]. If we treat the molecule as a collection of spherical atoms, using Eq. (4) the total electrostatic solvation energy is
@@
where $r_{ij}$ is the distance between the atomic centers of atoms $i$ and $j$. Here the first equation of $E_{\text{solv}}$ has two terms: a sum of individual Born’s terms and pairwise Coulombic terms. Note the actual molecule is better represented by a dielectric interface (e.g. solvent excluded surface (SES)) which separates inside domain $\Omega_{1}$ and outside domain $\Omega_{2}$. The charge $q_{i}$ is located at the center ${\bf r}_{i}$ of a sphere with radius $a_{i}$. We assume $\Omega_{1}$ contains all these spheres. The model using dielectric interface requires a step further as been approximated in the second equation in Eq. (5).
Here $f_{ij}^{\text{GB}}$ is the effective Born radii ($i=j$) or effective interaction distance ($i\neq j$). Assuming the effective Born Radii $R_{i}$’s are obtained, a popular estimation of $f_{ij}^{\text{GB}}$ is:
$f_{ij}^{\text{GB}}(r_{ij})=(r_{ij}^{2}+R_{i}R_{j}e^{\frac{-r_{ij}^{2}}{4R_{i}R_{j}}})^{\frac{1}{2}}$
Note $R_{i}$ depends not only on $a_{i}$, but also on radii and relative positions of all other atoms. The estimation of effective Born radii is an active research area [52]. From Poisson-Boltzmann theory, the perfect Born radii is given as
$R_{i}=\frac{1}{2}(\frac{1}{\epsilon_{2}}-\frac{1}{\epsilon_{1}})\frac{q_{i}}{E^{\text{PB}}_{\text{solv},i}}$
with $E^{\text{PB}}_{\text{solv},i}$ as the electrostatic solvation energy from the PB model with all except $i$th charge muted. This PB-GB relationship can be seen in Fig. 1(c), in which the electrostatic solvation energy of the sphere equals that of the molecule, both with a charge at the atomic center. There are many approaches to approximate $R_{i}$ and a typical one is the volume integration via quadrature as
$R_{i}^{-1}=\frac{1}{4\pi}\int_{\Omega_{2}}\frac{q_{i}^{2}}{r^{4}}d{\bf r}=\frac{q_{i}^{2}}{a_{i}}-\frac{1}{4\pi}\int_{\Omega_{1}/B({\bf r}_{i},a_{i})}\frac{q_{i}^{2}}{r^{4}}d{\bf r}.$

### v3

@@ Abstract
In this project, we present a deep neural network (DNN)-based biophysics model that uses multi-scale and uniform topological and electrostatic features to predict protein properties, such as Coulomb energies or solvation energies. The topological features are generated using element-specific persistent homology (ESPH) on a selection of heavy atoms or carbon atoms. The electrostatic features are generated using a novel Cartesian treecode, which adds underlying electrostatic interactions to further improve the model prediction. These features are uniform in number for proteins of varying sizes; therefore, the widely available protein structure databases can be used to train the network. These features are also multi-scale, allowing users to balance resolution and computational cost. The optimal model trained on more than 17,000 proteins for predicting Coulomb energy achieves MSE of approximately 0.024, MAPE of 0.073 and $R^{2}$ of 0.976. Meanwhile, the optimal model trained on more than 4,000 proteins for predicting solvation energy achieves MSE of approximately 0.064, MAPE of 0.081, and $R^{2}$ of 0.926, showing the efficiency and fidelity of these features in representing the protein structure and force field. The feature generation algorithms also have the potential to serve as general tools for assisting machine learning based prediction of protein properties and functions.
[…]
{tocentry}[Uncaptioned image]
@@ 1 Introduction
One of the overarching themes of biology is that structure determines function, which becomes more significant when we explore the biological world at the molecular level. With the advancement and availability of structure determination techniques such as X-ray crystallography Bragg:1912, Crowfoot:1935, NMR spectroscopy Aue:1976, cryo-electron microscopy Cheng:2015, etc., there are rapidly growing 3D structures of protein, nucleic acids, and complex assemblies, etc. on publicly accessible repository such as Protein Data Bank (PDB) Berman:2000, GenBank Benson:2012, etc. These structure determination techniques have strengths and limitations that make them suitable for different types research studies. For examples, X-ray crystallography, due to the requirement of crystallization, is best for static high-resolution structures. NMR is valuable for studying dynamics in solution, particularly for the studies of smaller proteins and complexes. Cryo-electron microscopy which becomes more and more popular recently, excels in visualizing large, complex, and flexible biomolecular systems.
@@
In addition to experiments, computing approaches are also available for the prediction of protein structures and properties. The main stream of Machine Learning (ML) based protein prediction focuses on using amino acid sequence to predict protein structure or structure-related properties, resulting in the trending protein language models (PLM) Bepler:2021, Weissenow:2025, which is a transformer-based model inspired by Natural Language Processing (NLP). PLM has a wide range of applications such as protein pKa prediction Xu:2025, protein structure prediction Lin:2023, protein design Madani:2023, protein evolution Zhang:2024, etc. Recently, the appearance of AlphaFold developed by Google DeepMind revolutionarily reshaped the structures determination techniques Abramson:2024.
@@
Both PLM and AlphaFold take protein sequences as inputs to predict structures or embeddings which keep important features. However, the development of ML methods in protein property predictions is hindered by the difficulty to represent protein structures and force fields using reliable features. To this end, the structure data, which is largely variant from protein to protein, needs to be converted to uniform size of features to be loaded into the ML models. Successful cases are the use of graph convolutional network (GCN) for predicting protein functions by features extracted from a protein language model and protein structure Gligorijevic:2021, protein-DNA binding specificity prediction using a wide range of features from geometry Sagendorf:2024, Mitra:2025, protein-ligand binding affinity prediction using topological features, Wang:2017, Wee:2022, molecular structure representation using persistent diagram Townsend:2020, etc. However, the important electrostatic interactions are often ignored in the formation of the protein features because of the challenges from long-range and pairwise nature of the electrostatic interactions.
@@
In this paper, we provide a novel approach that utilizes the far-field expansion in the Cartesian treecode algorithm Li:2009 to generate electrostatic features. To this end, the charges at the atomic centers are represented by multipole moments at the cluster centers, which inherit the key idea of the treecode algorithm Li:2009: replacing pairwise particle-particle interactions with particle-cluster interaction. By combining electrostatic features with the topological features generated using a modification of the persistent homology approach used by Cang and Wei Cang:2017, we represent the protein structures and the associated force field in a uniform and multiscale fashion. With these features, we train a deep neural network (DNN) to predict protein Coulombic energy and electrostatic solvation energy efficiently and accurately. Here, the Coulomb Energy and Solvation energy used as labels are generated by pair-wise Coulombic interactions and the surrogate model for solving the Poisson-Boltzmann (PB) model using a second-order Matched Interface and Boundary (MIB) method based PB solver Chen:2011. Apart from the recent research using Physics Informed Neural Network (PINN) to solve the Poisson-Boltzmann model Achondo:2025, Chen:2025, the present work is fundamentally different as a data-driven method with uniform and multiscale features.
There exist machine learning methods for the prediction of solvation energy under the framework of the PB model. For examples, Huang et al. developed a neural-network formulation for solving the dielectric-boundary PB equation by minimizing a variational electrostatic free-energy functional, enabling direct prediction of PB electrostatic solvation energies while preserving the physical structure of the continuum model Huang:2026; Röcken et al. proposed the ReSolv framework trained using experimental hydration free energies and ab initio data, achieving near experimental accuracy while accelerating solvation energy prediction Rocken:2024; Ferraz-Caetano et al. introduced an explainable supervised learning approach based on ensemble regression models that predicts solvation energies from physicochemical descriptors FerrazCaetano:2024; Wu and Luo presented SPACIER Wu:2025, a large-scale benchmark dataset of solvation energies grounded computed by using AMBER-PBSA Case2023AmberTools. The major advantage of the proposed method is at the feature generation independent of the PB framework, thus in addition to this particular applications in predicting Coulombic energy and PB-based electrostatic solvation energy, the feature generation algorithm developed in this work potentially has broader applications to many other ML models, as far as protein structures and force fields are the inputs represented uniformly across proteins of varying sizes.
@@
This manuscript is organized as follows. Following this introduction, we present the methods section, which includes both the modeling framework and numerical algorithms, followed by the simulation results. The manuscript ends with a section of conclusion.
2 Methods
@@
The objective of this project is to design uniform and multiscale features that represent protein structure and electrostatics, and then to use ML models, such as DNN, to evaluate the effectiveness of these features. The models are trained using labels such as solvation energy computed from surrogate models, e.g. the numerically solved Poisson-Boltzmann model. We hereby introduce the models and algorithms involved in this work. To study protein structure–property relationships, we focus on the atomic description and the electrostatic interactions. The protein properties considered in this paper are Coulomb energy $E_{\text{coul}}$ and electrostatic solvation energy $E_{\text{solv}}$. For a protein, the Coulomb energy $E_{\text{coul}}$ is the energy required to move atomic charges from infinity to their present positions(i.e. atomic centers specified in the Protein Data Bank (PDB) file. The electrostatic solvation energy $E_{\text{solv}}$ is the electrostatic free energy required to solvate the protein in a solvent, such as water. Formal quantitative definitions of these two properties are provided through mathematical models. For the protein-solvent interaction, we use the implicit solvent model in which water is treated as a continuum in order to reduce computational cost. From a comparative perspective, we introduce the two most popular implicit solvation models, namely the Poisson-Boltzmann model and the Generalized Born’s model, which are used to generate labels or features. We then introduce persistent homology and describe how it is used to construct topological features. Finally, we present the Cartesian treecode Li:2009 algorithm, and its modification into a point-multipole representation, which forms the basis for generating electrostatic features.
In this project, we adopt a combination of electrostatic features and topological features for the following reasons. Electrostatic features provide a uniform and multiscale representation of the atomic charges, accounting for both charge quantities and their locations. Topological features, on the other hand, encode information derived from the atomic centers of selected atoms, such as alpha-carbon atoms or heavy atoms. These features reveal the intrinsic topological invariants, which are not directly represented by electrostatic features. Together, these two types of features efficiently represent charge distributions, their locations, and underlying topological information.
@@ 2.1 The Poisson-Boltzmann model
Figure 1 depicts the popular implicit solvent models. In Fig.  1(a) a protein is represented by a collection of $N_{c}$ spherical atoms with centered partial charges. The molecular surface $\Gamma$ (also known as the solvent excluded surface Richards:1977, Connolly:1985) is defined by the trace of a water molecule represented by a red sphere rolling on contacting with the protein atoms. The Poisson-Boltzmann model is shown in Fig. 1(b), where the molecular surface $\Gamma$ divides the entire computational domain $\Omega$ into the protein domain $\Omega_{1}$ with dielectric constant $\varepsilon_{1}$ and atomic charges $q_{k}$ located at ${\bf x}_{k},k=1\,{:}\,N_{c}$, and the solvent domain $\Omega_{2}$ with dielectric constant $\varepsilon_{2}$ and dissolved mobile ions. Assuming a Boltzmann distribution for the ion concentration, and considering the case of two ion species with equal and opposite charges (e.g. Na+, Cl-), in the limit of weak electrostatic potential one obtains the linearized Poisson-Boltzmann (PB) model Baker:2005,
$\varepsilon_{1}{\rm NewA}^{2}\phi_{1}(\textbf{x})$ $=-{4\pi}\sum_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k}),\quad{\bf x}\in\Omega_{1},$ (1a)
$\varepsilon_{2}{\rm NewA}^{2}\phi_{2}(\textbf{x})$ $=\kappa^{2}\phi_{2}(\textbf{x}),\quad{\bf x}\in\Omega_{2},$ (1b)
@@
Figure 1: Implicit solvent models. (a) The Molecular Surface $\Gamma$: the trace of the solvent probe (shown in red) when it is rolled on contacting the spherical atoms of the protein (shown in green with a centered charge); (b) The Poisson–Boltzmann Model: two domains $\Omega_{1}$ (dielectric constant $\epsilon_{1}$ with partial charges as weighted summation of delta functions) and $\Omega_{2}$ (dielectric constant $\epsilon_{2}$ with mobile ions modeled by the Boltzmann distribution) are separated by the molecular surface $\Gamma$; (3) The Generalized Born Model: the protein is represented as a collection of $N_{c}$ spherical atoms centered at ${\bf x}_{k}$ with charge $q_{k}$ and Born Radius $R_{k}$ (only the $k$th atom is shown).
The PB model governs the electrostatic potential $\phi$ in the entire space. Theoretically after $\phi$ is obtained, its gradient will produce electrostatic field while its integral will generate potential energy. However, there are many challenging issues on properly obtaining the field and energy (e.g. definition of field on molecular surface $\Gamma$ Gilson:1993, Geng:2011). Our attention for this project is on the energy side as described below. The electrostatic free energy is given as
@@
where $\rho({\bf x})=\sum\limits_{k=1}^{N_{c}}q_{k}\delta(\textbf{x}-\textbf{x}_{k})$ is the charge density as a sum of partial charges weighted delta functions and the $E_{\text{solv}}=\frac{1}{2}\sum\limits_{k=1}^{N_{c}}q_{k}\phi_{\text{reac}}({\bf x}_{k})$ term is the electrostatic solvation energy. The $\phi_{\text{reac}}({\bf x}_{k})$ is the reaction potential at ${\bf x}_{k}$ as the remaining component when Coulomb potential $\phi_{\text{coul}}({\bf x}_{k})=\sum\limits_{j=1,j\neq k}^{N_{c}}\frac{q_{j}}{\epsilon_{1}|{\bf x}_{k}-{\bf x}_{j}|}$ is taken away from the electrostatic potential $\phi({\bf x}_{k})$.
Solving the PB model numerically by grid-based methods is challenging because (1) the protein is represented by singular point charges, (2) the molecular surface is geometrically complex, (3) the dielectric constant is discontinuous across the surface, (4) the domain is unbounded. To overcome these numerical difficulties, many finite difference interface methods have been developed Chern:2003, LeVeque:1994, Qiao:2006. For the purpose of dealing with arbitrarily shaped dielectric interfaces based on a simple Cartesian grid, a matched interface and boundary (MIB) PB solver Zhou:2006a, Yu:2007, Geng:2007, Chen:2011, Geng:2017 has been developed through rigorous treatments of geometrical and charge singularities. Alternatively, boundary element methods (BEM) for the PB model were developed later Zauhar:1988, Juffer:1991, Liang:1997, Boschitsch:2002, Bordner:2003, Lu:2007, Geng:2013b, Chen:2023 with several inherent advantages, (1) only the molecular surface is discretized rather than the entire solute/solvent volume, (2) the atomic charges are treated analytically, (3) the interface conditions are accurately enforced, (4) the far-field boundary condition is imposed analytically, (5) the matrix-vector product in iterative methods can be accelerated by fast algorithms Barnes:1986, Greengard:1987, Li:2009, Tausch:2003. In this paper, we computed the solvation energies using the MIBPB solver Chen:2011 since accuracy is of critical importance for the generation of labels ($E_{\text{solv}}$) that will be used to train the neural network.
@@ 2.2 Generalized Born’s Model
If ion effects are not considered, the solvation energy and the reaction potential can also be efficiently computed by using the generalized Born’s (GB) model, an approximation to the Poisson’s model. The GB model has significantly reduced computational cost as opposed to the PB or Poisson models, whose solutions require solving a 3D partial differential equation. Although we did not use the GB model to produce any numerical results for this project, we describe the GB model for its relationship with the PB model and its potential usage in electrostatic interactions. In fact, the GB model can calculate the electrostatic solvation energy which can be used either as the label or as a feature Chen:2021, Chen:2024 in our ML framework. It can also produce the reaction potentials at the atomic center, which can be used to generate the second level of electrostatic features as shown in Fig. 4 to be explained within that content.
Under the GB model, the total electrostatic solvation energy is
@@
where $r_{ij}$ is the distance between the atomic centers of atoms $i$ and $j$, $f_{ij}^{\text{GB}}$ is the effective Born radii ($i=j$) or effective interaction distance ($i\neq j$). This PB-GB relationship can be seen in Fig. 1(c), in which the electrostatic solvation energy of the sphere equals that of the molecule, both with a charge at the atomic center. Some details about the GB model can be found in the supplement material.